x-intercept and slope equation form

Grade:12th pass

1 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear amarnath,

We know that slope-Intercept form of line is

Y=mX+c

Here,

m = slope of the line in xy plane

c = Y intercept if the line

if you write the same equation in another form(putting X on LHS)

X = (1/m)Y - c/m

if an equaiton is in this form

m = slope of the line in xy plane

-c/m = X intercept of the line

GOODLUCK

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum..